OpenStudy (freemap):

What is the length of the segment joining the points at (–1, 6) and (–5, 3)?

OpenStudy (freemap):

I got 4.47

OpenStudy (photon336):

you could use the distance formula. \[\sqrt{(x_{2}-x_{1})^2+(y_{2}-y_{1})^2}\]

OpenStudy (popitree):

its 5

OpenStudy (freemap):

6--1*2 +3--5*5 6^2 +8^ 36+64 100 square root of 100 is 10

OpenStudy (freemap):

oh o just seen your answer @popitree

OpenStudy (freemap):

what did i do wrong?

OpenStudy (popitree):

-1-(-5) = -4 (-4)^2 = 16 6-3 = 3 3^2 = 9 sqrt(16+9) = 5

OpenStudy (popitree):

hope u got it

OpenStudy (freemap):

oh, I thought you start off with x2 which is 6 I'm sorry I just want to make sure I understand

OpenStudy (jdoe0001):

yea, same as the one before

OpenStudy (photon336):

\[(-1-(-5))^2+(6-3)^2 = 16+9 = \sqrt{25} = 5\]

OpenStudy (photon336):

I agree with @popitree

OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ -1}}\quad ,&{\color{blue}{ 6}})\quad % (c,d) &({\color{red}{ -5}}\quad ,&{\color{blue}{ 3}}) \end{array}\qquad % distance value d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)

OpenStudy (freemap):

Okay I get it instead of saying x2 -x 1 it would be x1 - x2

OpenStudy (freemap):

Thanks @popitree @Photon336 and @jdoe0001